OpenStudy (anonymous):

Graph the linear function using the slope and the y-intercept:f(x)=-2x+7

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/ejewatoji8

OpenStudy (anonymous):

@SolomonZelman This did not answer my question

OpenStudy (solomonzelman):

okay, lets plug in some x-values \(\normalsize\color{blue}{ f(x)=-2x+7}\) When x=0 \(\normalsize\color{blue}{ f(x)=-2(0)+7~~~~⇒~~~~f(x)=7}\) Thus we get (0,7) (this is the y-intercept) When x=1 \(\normalsize\color{blue}{ f(x)=-2(1)+7~~~~⇒~~~~f(x)=5}\) Thus we get (1,5) Plot the points and connect them with the line.

OpenStudy (solomonzelman):

I am not a very good teach, in fact an awful teacher-:(

OpenStudy (anonymous):

So how would i Graph that?

OpenStudy (anonymous):

https://www.youtube.com/watch?v=rFaFBoWnD8Q

OpenStudy (anonymous):

@SolomonZelman